# Federated Learning over Connected Modes

### v2

@@ 4.2 Results
SimpleCNN pre-trained ResNet-18 SimpleCNN pre-trained
@@
FedAvg 66.65 67.08 68.22 68.74 67.35 67.74 76.17 76.49 77.47 79.75 76.80 78.79
FedProx 66.14 66.80 67.79 67.67 67.88 68.00 75.72 75.88 76.91 78.21 76.39 78.06
Ditto 66.77 79.69 67.96 74.46 69.01 83.29 75.84 80.53 76.92 80.05 76.90 83.12
Floco 68.26 80.92 69.79 74.64 74.61 87.38 79.11 82.29 77.95 83.89 77.14 83.54
@@
SimpleCNN pre-trained ResNet-18 SimpleCNN pre-trained
@@
FedAvg 14.73 17.03 14.08 16.61 5.50 13.02 2.12 10.20 14.02 17.11 13.80 17.33
FedProx 14.73 16.48 13.64 16.67 5.00 12.60 1.93 10.19 13.47 17.70 13.39 17.71
Ditto 14.68 11.56 14.23 14.18 5.00 9.73 1.96 9.66 14.20 16.38 13.74 13.93
Floco 10.98 9.84 10.12 12.95 2.68 7.62 1.91 9.21 12.96 13.95 13.28 13.33
Table 3: Average local test accuracy for the 5 % worst performing clients.
CIFAR-10 FEMNIST
SimpleCNN pre-trained ResNet-18 SimpleCNN pre-trained
5-Fold Dir(0.5) 5-Fold Dir(0.5) SqueezeNet
FedAvg 0.42 $\pm$ 0.03 0.51 $\pm$ 0.04 0.50 $\pm$ 0.06 0.65 $\pm$ 0.03 0.46 $\pm$ 0.08 0.42 $\pm$ 0.12
FedProx 0.45 $\pm$ 0.05 0.51 $\pm$ 0.04 0.52 $\pm$ 0.02 0.64 $\pm$ 0.03 0.44 $\pm$ 0.10 0.42 $\pm$ 0.09
Ditto 0.65 $\pm$ 0.02 0.60 $\pm$ 0.03 0.72 $\pm$ 0.04 0.67 $\pm$ 0.05 0.37 $\pm$ 0.20 0.45 $\pm$ 0.19
Floco 0.66 $\pm$ 0.03 0.59 $\pm$ 0.03 0.79 $\pm$ 0.02 0.72 $\pm$ 0.03 0.55 $\pm$ 0.07 0.53 $\pm$ 0.13
Table 3 and 3 summarize the main experimental results, where Floco consistently outperforms the baselines across the different experiments in terms of global (red) and local (blue) test accuracy, as well as test ECE. The global and local test metrics are measured after the last communication round and averaged over 5 different seed runs. Below we report on detailed observations.
@@
We first evaluate the global and local test performance on CIFAR-10 with the non-IID data splits generated by the 5-Fold and $\mathrm{Dir}(\beta)$ procedures, as well as the natural non-IID data splits in the FEMNIST dataset. Table 3 shows the test accuracies on CIFAR-10 with SimpleCNN trained from random initialization (left) and ResNet-18 fine-tuned from the ImageNet pre-trained model (middle), respectively. It also shows the test accuracies on FEMNIST with SimpleCNN trained from random initialization and SqueezeNet fine-tuned from the ImageNet pre-trained model (right). We clearly see that Floco outperforms all baselines in terms of global (red) as well as average local (blue) test accuracy.
In some cases, e.g., SimpleCNN on the CIFAR-10 Dir(0.5) split, as well as pre-trained SqueezeNet on the FEMNIST, the performance gain from the Ditto baseline is marginal. However, Floco is still beneficial, because training Ditto requires roughly $2$ times larger computation and storage costs than Floco, since the overhead added to FedAvg by Floco is minimal, see Section 4.3. Furthermore, the personalization procedure of Ditto can be straightforwardly integrated into the Floco framework, boosting its performance and significantly outperforming Ditto (We call this method Floco+, which is introduced and evaluated in Appendix C). Besides test accuracy, we also observe TTA improvement of our method for all settings of up to ×$\times 4$ for average local test accuracy and ×$\times 3.8$ for global test accuracy than the next best baselines. We report all TTAs in Table 4.
@@
Next, we evaluate and benchmark the quality of uncertainty estimation of all methods. For this purpose we evaluate the ECE on each model-dataset combination for each baselines and show the results in Table 3. As shown, Floco yields better ECE for every setting, ranging from 1.23-3.52 and 0.05-3.52 of local and global ECE improvement compared to the next best baseline, respectively.
@@
Similar to Table 3, we plot the TTA improvement for Floco. In particular, we show the TTA improvement of Floco over FedAvg and Ditto. Again, all reported results are average values that we computed over 5 different random seeds.
@@
SimpleCNN pre-trained ResNet-18 SimpleCNN pre-trained
@@
Floco vs. FedAvg x1.5 x4.4 x1.6 x3.7 x3.5 x4.33 x1.8 x2.6 x1.4 x2 x1.1 x2.4
Floco vs. Ditto x1.5 x1.2 x1.3 x1.8 x3.8 x1.9 x2.7 x1.9 x2.4 x1.9 x1 x1.1
@@ 4.3 Analysis and Discussion
Here we provide additional analyses and discussion on Floco.
@@
Figure 2 shows the test accuracy curves during training for global (left) and local (center) test accuracies of different methods with the standard deviation over 5 trials as shadows. We observe that Floco not only converges to a better global and local test accuracy, but also shows small standard deviation across trials. The latter implies that our systematic regularization through the solution simplex stabilizes the training dynamics significantly. Figure 2 (right) shows the total gradient variance—the sum of the variances of the updates $\Delta\boldsymbol{w}^{t}_{k}=\boldsymbol{w}^{t}_{k}-\boldsymbol{w}_{0}^{t-1}$ for FedAvg, and $\Delta\boldsymbol{\theta}^{t}_{m,k}=\boldsymbol{\theta}^{t}_{m,k}-\boldsymbol{\theta}_{m,0}^{t-1}$ for Floco, respectively. More specifically, we compute the variance over the last fully-connected layer, given by
$\sigma_{\mathrm{total}}^{2}(t)$ $=\textstyle\sum_{k\in\mathcal{S}^{t}}\|\Delta\boldsymbol{w}^{t}_{k}-\frac{1}{|\mathcal{S}^{t}|}\sum_{k\in\mathcal{S}^{t}}\Delta\boldsymbol{w}^{t}_{k}\|_{2}^{2},$ (12)
@@
$\sigma_{\mathrm{total}}^{2}(t)$ $=\textstyle\sum_{k\in\mathcal{S}^{t}}\|\Delta\overline{\boldsymbol{\theta}}_{k}^{t}-\frac{1}{|\mathcal{S}^{t}|}\sum_{k\in\mathcal{S}^{t}}\Delta\overline{\boldsymbol{\theta}}_{k}^{t}\|_{2}^{2},$ (13)
with $\Delta\overline{\boldsymbol{\theta}}_{k}^{t}=\frac{1}{M+1}\sum_{m=1}^{M+1}\Delta\boldsymbol{\theta}_{m,k}^{t}$, for Floco. As discussed in [39, 40], a small total variance indicates effective collaborations with consistent gradient signals between the clients, leading to better performance. From the figure, we see that the total gradient variance of Floco is much lower and more stable than the baseline methods, which, together with its good performance observed in Table 3, is consistent with their discussion. The variance reduction with Floco implies that the degrees of freedom of the solution simplex can absorb the heterogeneity of clients to some extent, making the gradient signals more homogeneous. Moreover, [41] argued that the last classification layer has the biggest impact on performance, implying that reducing the total variance of the classification layer, as Floco does with simplex learning, is most effective.
Figure 2: Global (left) and average local (center) test accuracy for SimpleCNN on FEMNIST. We can clearly observe a jump in average local test accuracy at $\tau=200$, which is a result of our subregion assignment. Right shows the total variance of the gradients for the last fully-connected layer.
@@ 5 Related Work
There are a few existing works that apply simplex learning to federated learning. [29] showed that enforcing a low loss simplex between independently initialized global and client models yields good personalized FL performance. Their approach builds on [19], which finds optimal interpolation coefficients between a global and local model to improve personalized FL. However, their simplex is restricted to be 1D, i.e., a line segment, and the global model performance is comparable to the plain FedAvg. Moreover, they train a solution simplex over all layers between global and local models, which is computationally expensive and limits its applicability to training from scratch. This should be avoided if pre-trained models are available [32, 43]. Our method generalizes to training low-loss simplices of higher dimensions in a FL setting, tackles both the global and personalized FL objectives, is applicable to pre-trained models, and shows significant performance gains by employing our proposed subregion assignment procedure.
@@ 6 Conclusion
FL on highly non-IID client data distributions remains a challenging problem and a very actively researched topic. Recent works tackle non-IID FL settings either through global or personalized FL. While the former aims to find a single optimal set of parameters that fit a global objective, the latter tries to optimize multiple local models each of which fits the local distributions well. These two different objectives may pose a trade-off, that is, personalized FL might adapt models to strongly to local distributions which might harm the global performance, while global FL solutions might fit none of the local distributions if the local distributions are diverse. In this paper, we addressed this issue by leveraging the mode-connectivity of neural networks. Specifically, we propose Floco, where each client trains an assigned subregion within the solution simplex, which allows for personalization, and at the same, contributes to learning a well-performing global model. Floco achieves state-of-the-art performance in both global and personalized FL, with minimal computational and communication overhead during training and no overhead during inference.
@@ Appendix A Optimization Problem
The Lagrangian of the lower-level optimization problem in (9) has the following formulation $\mathcal{L}(\boldsymbol{\alpha}_{k},\lambda)=\frac{1}{2}\|\boldsymbol{\alpha}_{k}-\boldsymbol{\kappa}_{k}\|_{2}^{2}+\lambda({\bf 1}^{T}\boldsymbol{\alpha}_{k}-z)$, with $\lambda\in\mathbb{R}$ being the Langrange multiplier. The Lagrangian can be further rewritten to $\mathcal{L}(\boldsymbol{\alpha}_{k},\lambda)=\frac{1}{2}\|\boldsymbol{\alpha}_{k}-(\boldsymbol{\kappa}_{k}-\lambda{\bf 1})\|_{2}^{2}+\lambda({\bf 1}^{T}\boldsymbol{\kappa}_{k}-z)-\lambda^{2}n$ such that the optimization problem reduces to solving:
@@ Appendix B Training Hyperparameters
Table 6 summarizes all hyperparameters that were used for each dataset/model combination. We train all models for a total of 500 communication rounds, except for ResNet-18 on CIFAR-10, which we train for 100 communication rounds. Moreover, we train each setting using a total of 100 clients, and for FEMNIST we select a randomly chosen subset of 100 total clients for each trial, of which we select 10 randomly to participate in training in each communication round, except for SimpleCNN on CIFAR-10 where we select 30 out of 100 clients to participate in each round. We evaluate all clients after every ten communication rounds. For CIFAR-10 we train a SimpleCNN with batch size 50 using SGD with a learning rate of 0.02, momentum of 0.5, and weight decay of $10^{-5}$, and a pre-trained ResNet-18 with learning rate of batch size 32, using SGD with a learning rate of 0.01, momentum of 0.9, and weight decay of $10^{-4}$. For FEMNIST we train a pre-trained SqueezeNet with batch size 32 using SGD with a learning rate of 0.001, momentum of 0.9, weight decay of $10^{-4}$, and a SimpleCNN with batch size 32, learning rate 0.1, momentum of 0, weight decay of 0. For FedProx we set the proximity hyperparameter to $\mu=0.01$ for SimpleCNN on CIFAR-10, $\mu=0.1$ for pre-trained ResNet-18 on CIFAR-10, SimpleCNN on FEMNIST, and pre-trained SqueezeNet on FEMNIST. For Ditto we set the local epochs to the same value as epochs for the global model $E_{\textsc{Ditto}}=E$. All training hyperparameters for CIFAR-10 and FEMNIST on a SimpleCNN where taken from [29], CIFAR-10 on a pre-trained ResNet-18 from and FEMNIST on pre-trained SqueezeNet from [32].
@@
CIFAR-10/SimpleCNN 500 100 30 50 5 0.02 0.5 $10^{-5}$ 0.01
CIFAR-10/ResNet-18 100 100 10 32 5 0.01 0.9 $10^{-4}$ 0.1
FEMNIST/SimpleCNN 500 100 10 32 5 0.1 0.0 0.0 0.1
FEMNIST/SqueezeNet 500 100 10 32 5 0.001 0.9 $10^{-4}$ 0.1
Appendix C Combining Floco and Ditto
We can further enhance the personalized FL performance of Floco by additionally training a local model as in Ditto [14]. In this extension, called Floco+, each client personalizes the global endpoints $\{\widehat{\boldsymbol{\theta}}_{m}^{0}=\boldsymbol{\theta}_{m}\}_{m=1}^{M}$ by local gradient descent to minimize the Ditto objective, i.e.,
In Figure 3, we present preliminary results for Floco+ using the same hyperparameters as for Floco ($M=6$, $\tau=100$, and $\rho=0.1$). We evaluate Floco+ for SimpleCNN on CIFAR-10 with Dir(0.5) split because, in this setting, Floco only slightly surpasses the average local accuracy of Ditto in our main results (Table 3). We observe that Floco+ outperforms both Floco and Ditto. Note that this improvement comes with additional computational cost.
Figure 3: Average local test accuracy for Floco+, Floco and baselines when training SimpleCNN on CIFAR-10 with Dir(0.5) split.

### v4

@@ 4.2 Results
CifarCNN pre-trained ResNet-18 FemnistCNN pre-trained
@@
FedAvg 60.36 60.38 60.74 60.78 75.33 76.94 68.59 59.27 78.83 79.84 75.13 75.51
FedProx 60.68 60.36 60.40 60.27 76.93 77.46 62.27 60.26 78.84 80.15 75.47 75.99
FedPer 40.23 65.42 33.90 67.86 68.64 84.06 50.84 85.05 50.76 73.83 64.03 74.43
APFL 60.56 60.33 60.55 60.65 53.25 46.46 50.97 44.57 4.95 4.98 38.21 58.86
Ditto 60.36 72.22 60.74 73.90 75.33 69.18 68.59 76.23 78.83 82.02 57.89 65.06
FedRoD 56.36 74.03 46.12 76.42 17.46 31.82 10.27 33.85 4.95 4.99 4.95 4.95
Floco 62.93 71.78 62.57 71.04 77.15 85.90 73.62 80.38 78.99 84.09 75.86 77.00
Floco+ 62.93 75.08 62.57 76.50 77.15 84.88 73.62 85.89 78.99 84.75 75.86 82.41
@@
CifarCNN pre-trained ResNet-18 FemnistCNN pre-trained
@@
FedAvg 24.08 25.61 22.95 24.51 13.77 19.57 13.48 19.57 12.40 16.86 15.54 20.43
FedProx 23.76 25.56 23.19 24.89 12.40 12.41 15.16 19.83 12.41 16.93 15.48 20.04
FedPer 47.75 28.22 56.39 25.70 19.73 11.19 38.48 10.88 38.44 21.68 28.28 22.31
APFL 23.30 25.01 22.19 23.91 28.39 33.39 20.02 26.01 4.95 4.98 7.6 15.82
Ditto 24.08 19.13 22.95 17.64 13.77 16.43 13.48 14.50 12.40 14.65 15.54 18.06
FedRoD 29.78 18.40 41.91 17.45 75.59 64.07 89.31 64.07 4.95 4.99 4.99 4.99
Floco 21.82 18.44 20.06 18.75 11.48 9.44 10.30 11.28 10.28 13.94 14.65 19.15
Floco+ 21.82 17.69 20.06 16.50 11.48 12.42 10.30 11.98 10.28 13.87 14.65 15.35
Table 2 and 2 summarize the main experimental results, where Floco and Floco+ consistently outperform the baselines across the different experiments in terms of global (red) and local (blue) test accuracy, as well as test ECE. The global and local test metrics are measured after the last communication round and averaged over 5 different seed runs. The best performances are highlighted in bold, while the underlined entries indicate the settings that did not converge properly. Note that the global test performances of FedAvg and Ditto, as well as Floco and Floco+, are the same since they use the same global model. Below we report on detailed observations.
@@
We first evaluate the global and local test performance on CIFAR-10 with the non-IID data splits generated by the 5-Fold and $\mathrm{Dir}(\beta)$ procedures, as well as the natural non-IID data splits in the FEMNIST dataset. Table 2 shows the test accuracies on CIFAR-10 with CifarCNN trained from random initialization (left) and ResNet-18 fine-tuned from the ImageNet pre-trained model (center), respectively. It also shows the test accuracies on FEMNIST with FemnistCNN trained from random initialization (left) and SqueezeNet fine-tuned from the ImageNet pre-trained model (right). We clearly see that Floco and Floco+ outperform all baselines in terms of average local (blue) test accuracy by up to $6\%$, as well as global (red) by up to $5\%$.
@@
We evaluate and benchmark the quality of uncertainty estimation of all methods. For this purpose we evaluate the global as well as average local ECE on each model-dataset combination for each baseline on the test dataset and show the results in Table 2. As shown, Floco and Floco+ achieve better Expected Calibration Error (ECE) across all settings, with two exceptions: training a pre-trained ResNet-18 on the CIFAR-10 Dir(0.3) split and a pre-trained SqueezeNetV1 on FEMNIST. In the first case, the average local ECE for Floco and Floco+ is slightly worse than that of FedPer, suggesting mild overconfident for some clients. In the second case, the next best method (APFL) yields a significantly lower global test accuracy than our method, making a fair comparison of their ECE difficult.
@@
[…]
CIFAR-10 (CifarCNN)
5-Fold Dir(0.3)
FedAvg 44.0 $\pm$ 0.02 42.93 $\pm$ 0.03
FedProx 43.87 $\pm$ 0.02 43.23 $\pm$ 0.03
FedPer 52.67 $\pm$ 0.02 51.01 $\pm$ 0.02
APFL 43.27 $\pm$ 0.02 46.36 $\pm$ 0.03
Ditto 58.20 $\pm$ 0.03 58.69 $\pm$ 0.03
FedRoD 60.20 $\pm$ 0.02 61.12 $\pm$ 0.03
Floco+ 61.73 $\pm$ 0.02 61.13 $\pm$ 0.03
@@
Similar to Table 2, we plot the TTA improvement for Floco. In particular, we show the TTA improvement of Floco over FedAvg and FedProx, and the TTA improvement of Floco+ over Ditto, FedPer and FedRod, as all these methods include local fine-tuning. We report all TTAs in Table 4. The underlined entries indicate the cases where the test accuracies of our methods exceed the baseline method’s maximum accuracy already at the initial evaluation round, while the entries labeled ’x1.0’ represent the instances where our methods take the same evaluation rounds to achieve the baseline method’s maximum accuracy, i.e., comparable in terms of TTA. In addition to test accuracy, we also observe an improvement in Time-to-Accuracy (TTA) for our method across all settings.
@@
CifarCNN pre-trained ResNet-18 FemnistCNN pre-trained
@@
Floco vs. FedAvg x5.5 x4.6 x3.4 x3.1 x1.3 x1.8 x1.2 x8.0 x1.7 x1.2 x1.1 x1.1
Floco vs. FedProx x5.1 x4.9 x3.3 x3.8 x1.0 x1.8 x1.2 x9.0 x3.0 x1.2 x1.0 x1.1
Floco+ vs. Ditto x5.5 x2.3 x3.4 x2.1 x1.3 x2.0 x1.2 x1.7 x1.7 x4.0 x9.0 x4.0
Floco+ vs. FedPer x1.0 x1.5 x1.0 x1.3 x1.6 x1.5 x1.5 x1.5 x7 x7 x7.0 x2.7
Floco+ vs. FedRoD x9.4 x1.6 x24.5 x1.3 x10 x10 x10 x10 x7 x7 x10 x10
@@ 4.3 Analysis and Discussion
In this section, we provide further analyses and discussion on Floco.
@@
Figure 2 shows the test accuracy curves during training for global (left) and average local (center) test accuracies of different methods with the standard deviation over 5 trials as shadows. We observe that Floco and Floco+ not only converge faster than the global and pFL baselines respectively, but also show small standard deviation across trials. The latter implies that our systematic regularization through the solution simplex stabilizes the training dynamics significantly. Figure 2 (right) shows the total gradient variance—the sum of the variances of the updates $\Delta\boldsymbol{w}^{t}_{k}=\boldsymbol{w}^{t}_{k}-\boldsymbol{w}_{0}^{t-1}$ for FedAvg and FedProx (which almost overlap with each other), and $\Delta\boldsymbol{\theta}^{t}_{m,k}=\boldsymbol{\theta}^{t}_{m,k}-\boldsymbol{\theta}_{m,0}^{t-1}$ for Floco, respectively. More specifically, we compute the variance over the last fully-connected layer, given by
$\sigma_{\mathrm{total}}^{2}(t)$ $=\textstyle\sum_{k\in\mathcal{S}^{t}}\|\Delta\boldsymbol{w}^{t}_{k}-\frac{1}{|\mathcal{S}^{t}|}\sum_{k\in\mathcal{S}^{t}}\Delta\boldsymbol{w}^{t}_{k}\|_{2}^{2}$ (12)
@@
$\sigma_{\mathrm{total}}^{2}(t)$ $=\textstyle\frac{1}{M+1}\sum_{m=1}^{M+1}\sum_{k\in\mathcal{S}^{t}}\|\Delta\boldsymbol{\theta}_{m,k}^{t}-\frac{1}{|\mathcal{S}^{t}|}\sum_{k\in\mathcal{S}^{t}}\Delta\boldsymbol{\theta}_{m,k}^{t}\|_{2}^{2}.$ (13)
We have not plotted the gradient variances of Floco+ and the other pFL methods, since those are the same as for Floco and FedAvg, respectively. As discussed in [45, 46], a small total variance indicates effective collaborations with consistent gradient signals between the clients, leading to better performance. From the figure, we see that the total gradient variance of Floco is much lower and more stable, in terms of standard deviation, than the baseline methods, which, together with its good performance observed in Table 2, is consistent with their discussion. The variance reduction with Floco implies that the degrees of freedom of the solution simplex can absorb the heterogeneity of clients to some extent, making the gradient signals more homogeneous. Moreover, [47] argued that the last classification layer has the biggest impact on performance, implying that reducing the total variance of the classification layer, as Floco does with simplex learning, is most effective. As we show in the Appendix C, applying simplex learning to only the last layer, instead of learning a simplex in the whole parameter space, achieves faster personalized and global convergence.
Figure 2: Global (left) and average local (center) test accuracy for CifarCNN on CIFAR-10, 5-Fold. For Floco, we can clearly observe a jump in average local test accuracy at $\tau=250$, which is a result of our subregion assignment. Right shows the total variance of the gradients for the last fully-connected layer.
@@ 5 Related Work
There are few existing works that apply simplex learning to federated learning. [35] proposed SuPerFed, which enforces a low loss simplex between independently initialized global and client models, yielding good personalized FL performance. This approach builds on [25], which finds optimal interpolation coefficients between a global and local model to improve personalized FL. However, their simplex is restricted to be 1D, i.e., a line segment, and the global model performance is comparable to the plain FedAvg. Moreover, they train a solution simplex over all layers between global and local models, which is computationally expensive and limits its applicability to training from scratch. This should be avoided if pre-trained models are available [38, 49]. Our method generalizes to training low-loss simplices of higher dimensions in a FL setting, tackles both the global and personalized FL objectives, is applicable to pre-trained models, and shows significant performance gains by employing our proposed subregion assignment procedure. In Table 7 of Appendix E we benchmark Floco against the SuPerFed baseline on the CIFAR-10, 5-Fold, as well as Dir(0.5) splits using both a CifarCNN trained from scratch as well as a pre-trained ResNet18 on both global as well as local test performance, where we observe that Floco outperforms SuPerFed both in terms of global as well as local accuracy in all settings.
6 Limitations
In this work, we only evaluate our method on cross-silo FL settings with up to 100 clients. Unlike cross-device FL, which typically involves a much larger set of stateless clients (i.e., clients with limited data that hinders reliable modeling), our approach assumes stateful clients, each with sufficient data to enable effective grouping of similar clients. While our current analysis focuses on cross-silo FL, extending our method to the cross-device setting is an important direction for future research. Additionally, a thorough theoretical analysis of our approach remains a future research objective.
@@ 7 Conclusion
FL on highly non-IID client data distributions remains a challenging problem and a very actively researched topic. Recent works tackle non-IID FL settings either through global or personalized FL. While the former aims to find a single optimal set of parameters that fit a global objective, the latter tries to optimize multiple local models each of which fits the local distribution well. These two different objectives may pose a trade-off, that is, personalized FL might adapt models to strongly to local distributions which might harm the global performance, while global FL solutions might fit none of the local distributions if the local distributions are diverse. In this paper, we addressed this issue by leveraging the mode-connectivity of neural networks. Specifically, we propose Floco, where each client trains an assigned subregion within the solution simplex, which allows for personalization, and at the same, contributes to learning a well-performing global model. Floco achieves state-of-the-art performance in both global and personalized FL, with minimal computational and communication overhead during training and no overhead during inference.
@@
Acknowledgements
This work was funded by the German Ministry for Education and Research as BIFOLD - Berlin Institute for the Foundations of Learning and Data (ref. BIFOLD24B).
@@ Appendix A Optimization Problem
The Lagrangian of the lower-level optimization problem in (9) has the following formulation $\mathcal{L}(\boldsymbol{\alpha}_{k},\lambda)=\frac{1}{2}\|\boldsymbol{\alpha}_{k}-\boldsymbol{\kappa}_{k}\|_{2}^{2}+\lambda({\bf 1}^{T}\boldsymbol{\alpha}_{k}-z)$ with $\lambda\in\mathbb{R}$ being the Langrange multiplier. The Lagrangian can be further rewritten to $\mathcal{L}(\boldsymbol{\alpha}_{k},\lambda)=\frac{1}{2}\|\boldsymbol{\alpha}_{k}-(\boldsymbol{\kappa}_{k}-\lambda{\bf 1})\|_{2}^{2}+\lambda({\bf 1}^{T}\boldsymbol{\kappa}_{k}-z)-\lambda^{2}n$ such that the optimization problem reduces to solving
@@ Appendix B Training Hyperparameters
Table 6 summarizes all hyperparameters that were used for each dataset/model combination. We train CifarCNN on CIFAR-10 for a total of 500 communication rounds, ResNet-18 on CIFAR-10 for 100 communication rounds, FemnistCNN on FEMNIST for 350 rounds, and SqueezeNetV1 on FEMNIST for 1000 rounds. Moreover, we train each setting using a total of 100 clients, and for FEMNIST we select a randomly chosen subset of 100 total clients for each trial, of which we select 10 randomly to participate in training in each communication round, except for CifarCNN on CIFAR-10 where we select 30 out of 100 clients to participate in each round. We evaluate all clients after every ten communication rounds. For CIFAR-10 we train a CifarCNN with batch size 50 using SGD with a learning rate of 0.02, momentum of 0.5, and weight decay of $10^{-5}$, and a pre-trained ResNet-18 with learning rate of batch size 32, using SGD with a learning rate of 0.01, momentum of 0.9, and weight decay of $10^{-4}$. For FEMNIST we train a pre-trained SqueezeNet with batch size 32 using SGD with a learning rate of 0.005, momentum of 0, weight decay of $10^{-4}$, and a FemnistCNN with batch size 32, learning rate 0.1, momentum of 0, weight decay of 0. For FedProx we set the proximity hyperparameter to $\mu=0.01$ for all settings. For Ditto, FedRod and FedPer we set the local epochs to the same value as epochs for the global model, i.e. $E_{\textsc{Ditto}}=E$. All training hyperparameters for CIFAR-10 and FEMNIST on a FemnistCNN were taken from [35], CIFAR-10 on a pre-trained ResNet-18 from [49] and FEMNIST on pre-trained SqueezeNet from [38].
@@
CIFAR-10/CifarCNN 500 100 30 50 5 0.02 0.5 $10^{-5}$ 0.01
CIFAR-10/ResNet-18 100 100 10 32 5 0.01 0.9 $10^{-4}$ 0.01
FEMNIST/FemnistCNN 350 100 10 32 5 0.1 0.0 0.0 0.01
FEMNIST/SqueezeNetV1 1000 100 10 32 5 0.005 0.0 $10^{-4}$ 0.01
Appendix C Simplex Learning on all NN parameters
In Figure 5, we compare the global (left) and average client (right) test accuracy of Floco and Floco-All, where the latter applies simplex learning to all NN parameters. As expected, Floco-All converges to the same global and average local test accuracy, but needs more communication rounds to do so, since it needs to train more parameters.
Figure 3: Global test accuracy.
Figure 4: Average local client test accuracy.
Figure 5: Comparing simplex learning on all network layers vs. only on the last fully-connected layer.